# OceanSplat: Object-aware Gaussian Splatting with Trinocular View Consistency for Underwater Scene Reconstruction

### v1

@@ Abstract
We introduce OceanSplat, a novel 3D Gaussian Splatting-based approach for accurately representing 3D geometry in underwater scenes. To overcome multi-view inconsistencies caused by underwater optical degradation, our method enforces trinocular view consistency by rendering horizontally and vertically translated camera views relative to each input view and aligning them via inverse warping. Furthermore, these translated camera views are used to derive a synthetic epipolar depth prior through triangulation, which serves as a self-supervised depth regularizer. These geometric constraints facilitate the spatial optimization of 3D Gaussians and preserve scene structure in underwater environments. We also propose a depth-aware alpha adjustment that modulates the opacity of 3D Gaussians during early training based on their $z$-component and viewing direction, deterring the formation of medium-induced primitives. With our contributions, 3D Gaussians are disentangled from the scattering medium, enabling robust representation of object geometry and significantly reducing floating artifacts in reconstructed underwater scenes. Experiments on real-world underwater and simulated scenes demonstrate that OceanSplat substantially outperforms existing methods for both scene reconstruction and restoration in scattering media.
@@ Introduction
Underwater scene reconstruction is essential for several marine robotics tasks, including seafloor mapping (kapoutsis2016real; liu2020underwater; joshi2022underwater), ecological monitoring (marre2019monitoring; girdhar2023curee), and subsea infrastructure inspection (fang2023integration; huang2025visual). While visual data can generally support robotic operations effectively, the underwater optical properties, such as wavelength-dependent attenuation (akkaynak2017space), scattering (mcglamery1980computer; jaffe2002computer; chen2021underwater), and low illumination (marques2020l2uwe), significantly degrade perceptual cues (yu2023task; zhang2024dcgf) and hinder the deployment of vision-based autonomous systems (de2021impact). Consequently, achieving geometric understanding and spatial representation from this low-quality imagery is essential for practical operation of autonomous platforms. Although advances in remotely operated vehicles (ROVs) and autonomous underwater vehicles (AUVs) have greatly enabled the acquisition of large-scale underwater imagery (wynn2014autonomous), leveraging this data for scene representation remains highly challenging due to adverse visual conditions (beall20103d; huang2025visual).
@@
Figure 1: OceanSplat overcomes scattering and attenuation effects through trinocular view consistency, preserving object geometry and enabling high-quality underwater 3D reconstruction.
@@
To address these issues in underwater perception, the use of Neural Radiance Fields (NeRF) (mildenhall2020nerf) and 3D Gaussian Splatting (3DGS) (kerbl20233d), which have shown strong capability in 3D reconstruction, has emerged as a promising direction. However, these methods rely on assumptions tailored to in-air image acquisition, motivating recent research to integrate underwater image formation model (akkaynak2019sea).
@@
NeRF-based methods (sethuraman2023waternerf; zhang2023beyond; levy2023seathru; ramazzina2023scatternerf; tang2024neural) extend their volumetric rendering framework by integrating light-transmitting physics and scattering phenomena. However, these methods implicitly represent both object and backscatter components, which impedes accurate geometric representation and also results in slow rendering speed. More recently, 3DGS-based methods (yang2024seasplat; li2024watersplatting; wang2024uw) have been adapted for underwater environments, benefiting from their superior rendering performance and fast training. Despite notable progress in these studies, medium intensity is often absorbed into the 3D Gaussians, leading to numerous floating artifacts and degrading reconstruction quality.
@@
Unlike previous approaches, our goal is to robustly recover 3D geometry in underwater scenes subject to severe scattering and attenuation, without depending on external geometric supervision. To this end, we propose OceanSplat: Object-aware Gaussian Splatting for geometrically consistent underwater scene reconstruction. Our method enforces trinocular view consistency across translated camera views along two orthogonal axes, resolving multi-view inconsistencies and effectively guiding 3D Gaussians to align with the scene structure. Additionally, we derive a synthetic epipolar depth prior via triangulation across the translated views, and leverage it for self-supervised depth regularization. Furthermore, we apply a depth residual loss that constrains the $z$-component of each 3D Gaussian toward the alpha-blended depth to reduce floating artifacts. We also propose a depth-aware alpha adjustment that regulates the transparency of 3D Gaussians during the early stages of training based on their $z$-component and viewing direction, thereby discouraging the formation of medium-induced primitives. As shown in Figure 1, our method improves geometric accuracy for object representation under challenging scattering conditions by preventing the entanglement of 3D Gaussians with the scattering medium. Our key contributions are outlined in the following points:
@@
  • We propose OceanSplat, a novel underwater scene reconstruction method leveraging 3D Gaussian Splatting that employs trinocular view consistency to effectively preserve scene geometry in scattering media.
@@
  • We introduce a synthetic epipolar depth prior derived through triangulation from translated camera views, enabling self-supervised geometric regularization.
@@
  • Our approach surpasses contemporary state-of-the-art methods in both qualitative and quantitative evaluations across real-world and simulated datasets.
@@
Underwater scene reconstruction aims to represent precise scene structure and restore true color despite severe light attenuation and scattering. Initial approaches jointly estimated medium-dependent attenuation during reconstruction (skinner2017automatic) and exploited global context with active labeling for sonar-based modeling (debortoli2018real). Subsequent works (sethuraman2023waternerf; zhang2023beyond; tang2024neural) embed physics-based underwater light transport and scattering into NeRF (mildenhall2020nerf) to jointly recover geometry and true color appearance. Seathru-NeRF (levy2023seathru) leverages the SeaThru image formation model (akkaynak2019sea) within NeRF to successfully achieve novel view synthesis in real-world underwater environments. ScatterNeRF (ramazzina2023scatternerf) uses inverse rendering, while DecoNeRF (zhang2025decoupling) employs pseudo-labeling to disentangle scattering media from scene content and restore clear views in foggy scenes. Although these methods significantly improve reconstruction quality, they are often hindered by slow rendering speed and high memory consumption. Recently, 3DGS (kerbl20233d) has gained increasing attention for underwater scene reconstruction. SeaSplat (yang2024seasplat) incorporates underwater physics into 3D Gaussians to improve color restoration and structural fidelity, while WaterSplatting (li2024watersplatting) combines implicit medium modeling with explicit object representation for efficient, high-quality reconstruction. UW-GS (wang2024uw) further enhances geometry using pseudo-depth priors from foundation models (yang2024depth). Despite these advances, existing 3DGS-based methods struggle to represent geometry-consistent underwater scenes.
@@
Multi-view stereo (MVS) algorithms reconstruct detailed 3D geometry by leveraging images acquired from multiple calibrated viewpoints. Early approaches employed voxelization (kutulakos2000theory), dense point clouds (furukawa2009accurate), and depth maps (galliani2015massively; schonberger2016structure). In the deep learning era, MVSNet (yao2018mvsnet) introduced cost volumes to form geometry-aware 3D representations, greatly enhancing accuracy. Building upon these advancements, MVS has been integrated into neural rendering (mildenhall2020nerf; kerbl20233d), where NeRF-based methods (chen2021mvsnerf; wang2021neus; fu2022geo; wu2025sparis) utilize these concepts for surface reconstruction, while 3DGS-based approaches (chen2023neusg; guedon2024sugar; huang20242d; chung2024depth; chen2024pgsr; chen2024mvsplat; liu2024mvsgaussian; li2024dngaussian; huang2025fatesgs) harness multi-view relationships to enhance geometric fidelity. Among these, recent efforts (safadoust2024self; han2024binocular) have demonstrated promising results by constructing virtual stereo pairs to enforce binocular stereo consistency, effectively regularizing 3D Gaussians.
@@
3D Gaussian Splatting (kerbl20233d) models a scene with anisotropic 3D Gaussian primitives, each parameterized by a center position $\mu\in\mathbb{R}^{3}$, covariance matrix $\Sigma\in\mathbb{R}^{3\times 3}$, color $c$, opacity $\alpha$, and view-dependent appearance modeled by spherical harmonics. Each 3D Gaussian defines its spatial influence at a point $\mathbf{X}\in\mathbb{R}^{3}$ centered at $\mu$ as:
@@
For rendering, 3D Gaussians are projected onto 2D screen space and composited via alpha-blending of $N$ primitives that are sorted by depth and overlap at each pixel:
@@
To render a 3D underwater scene, the image formation is guided by the revised underwater image formation model (akkaynak2018revised), which decomposes the observed image into a direct component attenuated by the medium with respect to depth $z$, and a backscatter component accumulated along the viewing direction, as follows:
@@
Figure 2 shows the overall architecture of our method. We begin by initializing 3D Gaussians from a collection of RGB images $I\in\mathbb{R}^{H\times W\times 3}$, along with their corresponding camera intrinsics $K\in\mathbb{R}^{3\times 3}$ and extrinsic parameters $P\in\mathbb{R}^{4\times 4}$, which represent the homogeneous transformation matrix, recovered using Structure-from-Motion (SfM) (schonberger2016structure). We model medium properties along the ray $\vec{\mathbf{r}}$ using an MLP $\phi_{\text{med}}$, following (li2024watersplatting), which directly predicts attenuation $\sigma^{\text{attn}}$, backscattering $\sigma^{\text{bs}}$, and medium color $c^{\text{med}}$ as follows:
@@
We propose a trinocular stereo framework for 3DGS to improve object representation in scattering media. While prior works (safadoust2024self; han2024binocular) use only binocular stereo consistency, we show that a vertically translated viewpoint provides orthogonal constraints that complement horizontal stereo for 3D Gaussians. As shown in Figure 3, view-dependent sampling in alpha-blending often causes inconsistencies under minor viewpoint changes. Furthermore, previous stereo matching studies (okutomi1993multiple; imran2020unsupervised; shamsafar2021tristereonet) have shown that multi-baseline stereo surpasses single-baseline methods in accuracy and robustness. Inspired by these observations, we enforce trinocular view consistency with orthogonal baselines of varying lengths to better regularize the spatial position of 3D Gaussians.
@@
From a given camera pose $P_{c}$, we obtain the rasterized object and medium components, denoted by $I^{\text{obj}}_{c}$ and $I^{\text{med}}_{c}$, respectively. Two additional camera poses are defined by translating $P_{c}$ along the horizontal $h$ and vertical $v$ axes as:
@@
where $\mathbb{I}$ is the $3\times 3$ identity matrix, $\mathbf{t}_{h}=(b_{h},0,0)^{\top}$ and $\mathbf{t}_{v}=(0,b_{v},0)^{\top}$ are translation vectors defining the horizontally and vertically translated camera poses $P_{h}$ and $P_{v}$, respectively, with $b_{h},b_{v}\in\mathbb{R}$ denoting the baseline distances. Using these poses, we render two images $I^{\text{obj}}_{h}$ and $I^{\text{obj}}_{v}$ from the poses $P_{h}$ and $P_{v}$. We then apply disparity-based inverse warping to the rendered images. Let $D_{c}\in\mathbb{R}^{H\times W\times 1}$ be the rendered depth map from $P_{c}$. Based on stereo geometry, the horizontal and vertical disparity maps are computed as:
@@
where $f_{h},f_{v}\in\mathbb{R}$ denote the focal lengths in the horizontal and vertical directions. The resulting disparity maps $d_{h},d_{v}\in\mathbb{R}^{H\times W\times 1}$ represent the pixel-wise displacements needed to align the translated views with the central view, and $(x,y)$ denotes the spatial pixel coordinates. Then, the rendered images from the translated views, $I^{\text{obj}}_{h}$ and $I^{\text{obj}}_{v}$, are inverse-warped using disparity maps $d_{h}$ and $d_{v}$, respectively, in the direction opposite to the virtual camera shift, aligning them with the central image rendered from $P_{c}$. The resulting inverse warped images $I^{\text{obj}}_{h\rightarrow c},I^{\text{obj}}_{v\rightarrow c}\in\mathbb{R}^{H\times W\times 3}$ are defined as:
@@
Afterward, to emphasize dark regions as perceived in human vision, we follow prior studies (mildenhall2022nerf; levy2023seathru; li2024watersplatting), and employ a regularized L1 loss $L_{\text{R-L1}}$ defined as:
@@
To enhance geometric alignment, we introduce a synthetic epipolar depth prior $D_{\text{epi}}$, derived via triangulation between two translated views. For reliable correspondences in triangulation, we use 3D Gaussians sampled from the intersection of the view frusta of all trinocular camera views, with opacity above $\tau_{\alpha}$. Each selected Gaussian is projected onto views $P_{h}$ and $P_{v}$ via projection matrices $\mathbf{M}_{h},\mathbf{M}_{v}\in\mathbb{R}^{3\times 4}$, producing screen-space points $\mathbf{x}_{h}^{i},\mathbf{x}_{v}^{i}\in\mathbb{R}^{2}$. Let the homogeneous coordinates be $\tilde{\mathbf{x}}_{k}^{i}=[\mathbf{x}_{k}^{i\top},1]^{\top}\in\mathbb{R}^{3}$ and $\tilde{\mathbf{X}}_{i}=[\mathbf{X}_{i}^{\top},1]^{\top}\in\mathbb{R}^{4}$. We apply epipolar geometry to the projected correspondences, which yields a homogeneous linear system $\mathbf{A}_{i}\tilde{\mathbf{X}}_{i}=\mathbf{0}$, where $\mathbf{A}_{i}\in\mathbb{R}^{4\times 4}$ stacks the constraints from both views. To solve it stably, we rewrite it in the least-squares form:
@@
where $\mathbf{A}_{i}^{\prime}\in\mathbb{R}^{4\times 3}$ and $\mathbf{b}_{i}\in\mathbb{R}^{4}$ are obtained by separating the last column of $\mathbf{A}_{i}$. The triangulated point $\hat{\mathbf{X}}_{i}$ is transformed to the central camera view via extrinsics $\mathbf{R}_{c}\in\mathbb{R}^{3\times 3}$, $\mathbf{t}_{c}\in\mathbb{R}^{3}$, and the depth prior is defined as the $z$-component:
@@
After that, we apply an edge-aware Log-L1 loss (turkulainen2025dn) using $D_{\text{epi}}$ as follows:
@@
where $L_{\text{R-SSIM}}$ and $\lambda_{\text{s}}$ denote the Regularized SSIM loss (li2024watersplatting) and its corresponding weight, respectively, with $\lambda_{\text{s}}=0.2$. In addition to the baseline, we incorporate our proposed components into the final training objective:
@@
We evaluate underwater scene reconstruction on four scenes from SeaThru-NeRF dataset (levy2023seathru) and two scenes from In-the-Wild dataset (tang2024neural). In addition, to evaluate scene restoration performance, we generate simulated underwater and foggy scenes using the Fern scene from the LLFF dataset (mildenhall2019local). We simulate underwater scenes following previous methods (levy2023seathru) using normalized depth obtained from a depth foundation model (yang2024depth) with $\beta_{D}=[1.3,\ 1.2,\ 0.9]$, $\beta_{B}=[0.95,\ 0.85,\ 0.7]$, and $\beta_{\infty}=[0.07,\ 0.2,\ 0.39]$. For foggy scenes, we use $\beta_{D}=[0.5,\ 0.5,\ 0.5]$ and $\beta_{B}=\beta_{\infty}=1.2$. To validate our method, we compare scene quality using PSNR, SSIM (wang2004image), and LPIPS (zhang2018unreasonable).
@@
Our method builds on Nerfstudio (tancik2023nerfstudio), and employs a TCNN-based MLP (tiny-cuda-nn). We use 7K/3K training steps for densification/finetuning on SeaThru-NeRF (levy2023seathru) and simulated datasets, and 10K/5K steps on In-the-Wild (tang2024neural). At each iteration, $b_{v}$ is sampled from [0.4,0.4][-0.4,0.4], and $b_{h}$ is set to $1.5b_{v}$. We enable $L_{\text{tri}}$ and $L_{\text{epi}}$ during iterations 4K–8K and 6K–12K for 10K and 15K training steps, respectively. The $\tau_{\alpha}$ for synthetic epipolar depth is set to 0.8. The input resolution is increased from $\times 1/4$ to $\times 1/2$ at $1/5$ of total training steps, and from $\times 1/2$ to full resolution at $2/5$ to boost the training process. For depth-aware alpha adjustment, $w$ is set to 0.5 for $t<t_{\alpha}$ and is decayed to 0 at $t_{\alpha}$. We set the transition step $t_{\alpha}$ to 4K and 10K for 10K and 15K training steps, respectively. All experiments use a single RTX A6000.

### v2

@@ Abstract
We introduce OceanSplat, a novel 3D Gaussian Splatting-based approach for high-fidelity underwater scene reconstruction. To overcome multi-view inconsistencies caused by scattering media, we design a trinocular setup for each camera pose by rendering from horizontally and vertically translated virtual viewpoints, enforcing view consistency to facilitate spatial optimization of 3D Gaussians. Furthermore, we derive synthetic epipolar depth priors from the virtual viewpoints, which serve as self-supervised depth regularizers to compensate for the limited geometric cues in degraded underwater scenes. We also propose a depth-aware alpha adjustment that modulates the opacity of 3D Gaussians during early training based on their depth along the viewing direction, deterring the formation of medium-induced primitives. Our approach promotes the disentanglement of 3D Gaussians from the scattering medium through effective geometric constraints, enabling accurate representation of scene structure and significantly reducing floating artifacts. Experiments on real-world underwater and simulated scenes demonstrate that OceanSplat substantially outperforms existing methods for both scene reconstruction and restoration in scattering media.
@@ Introduction
Underwater scene reconstruction is essential for several marine robotics tasks, including seafloor mapping (Kapoutsis et al. 2016; Liu et al. 2020; Joshi et al. 2022), ecological monitoring (Marre et al. 2019; Girdhar et al. 2023), and subsea infrastructure inspection (Fang et al. 2023; Huang et al. 2025a). While visual data can generally support robotic operations effectively, the underwater optical properties, such as wavelength-dependent attenuation (Akkaynak et al. 2017), scattering (McGlamery 1980; Jaffe 2002; Chen et al. 2021b), and low illumination (Marques and Albu 2020), significantly degrade perceptual cues (Yu et al. 2023; Zhang et al. 2024) and hinder the deployment of vision-based autonomous systems (De and Pedersen 2021). Consequently, achieving geometric understanding and spatial representation from this low-quality imagery is essential for practical operation of autonomous platforms. Although advances in remotely operated vehicles (ROVs) and autonomous underwater vehicles (AUVs) have greatly enabled the acquisition of large-scale underwater imagery (Wynn et al. 2014), leveraging this data for scene representation remains highly challenging due to adverse visual conditions (Beall et al. 2010; Huang et al. 2025a).
@@
Figure 1: OceanSplat overcomes scattering and attenuation effects through trinocular view consistency, preserving object structure and enabling high-quality underwater 3D reconstruction.
@@
To address these issues in underwater perception, the use of Neural Radiance Fields (NeRF) (Mildenhall et al. 2020) and 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023), which have shown strong capability in 3D reconstruction, has emerged as a promising direction. However, these methods rely on assumptions tailored to in-air image acquisition, motivating recent research to integrate underwater image formation model (Akkaynak and Treibitz 2019).
@@
NeRF-based methods (Sethuraman et al. 2023; Zhang and Johnson-Roberson 2023; Levy et al. 2023; Ramazzina et al. 2023; Tang et al. 2024) extend their volumetric rendering framework by integrating light-transmitting physics and scattering phenomena. However, these methods implicitly represent both object and backscatter components, which impedes accurate geometric representation and also results in slow rendering speed. More recently, 3DGS-based methods (Yang et al. 2024a; Li et al. 2024a; Wang et al. 2024) have been adapted for underwater environments, benefiting from their superior rendering performance and fast training. Despite notable progress in these studies, medium intensity is often absorbed into the 3D Gaussians, leading to numerous floating artifacts and degrading reconstruction quality.
@@
Unlike previous approaches, our goal is to robustly recover scene geometry in underwater environments subject to scattering and attenuation, without depending on external geometric supervision. To this end, we propose OceanSplat: Object-aware Gaussian Splatting for geometrically consistent underwater scene reconstruction. Our method enforces trinocular view consistency across translated camera views along two orthogonal axes, resolving multi-view inconsistencies and effectively guiding 3D Gaussians to align with the scene structure. Additionally, we derive a synthetic epipolar depth prior through triangulation between the translated viewpoints and leverage it for self-supervised depth regularization. Furthermore, we apply a depth residual loss that constrains the $z$-component of each 3D Gaussian toward the alpha-blended depth to reduce floating artifacts. We also propose a depth-aware alpha adjustment that regulates the transparency of 3D Gaussians during the early stages of training based on their $z$-component and viewing direction, thereby discouraging the formation of medium-induced primitives. As shown in Figure 1, our method improves geometric accuracy for object representation under challenging scattering conditions by preventing the entanglement of 3D Gaussians with the scattering medium. Our key contributions are outlined in the following points:
@@
  • We propose OceanSplat, a novel underwater scene reconstruction method leveraging 3D Gaussian Splatting, which employs trinocular view consistency to align 3D Gaussians with scene geometry in scattering media.
@@
  • We introduce synthetic epipolar depth priors derived from translated virtual viewpoints via triangulation, enabling self-supervised depth regularization.
@@
  • Our approach surpasses existing state-of-the-art methods in both qualitative and quantitative evaluations across real-world and simulated datasets.
@@
Underwater scene reconstruction aims to represent precise scene structure and restore true color despite severe light attenuation and scattering. Initial approaches jointly estimated medium-dependent attenuation during reconstruction (Skinner et al. 2017) and exploited global context with active labeling for sonar-based modeling (DeBortoli et al. 2018). Subsequent works (Sethuraman et al. 2023; Zhang and Johnson-Roberson 2023; Tang et al. 2024) embed physics-based underwater light transport and scattering into NeRF (Mildenhall et al. 2020) to jointly recover geometry and true color appearance. Seathru-NeRF (Levy et al. 2023) leverages the SeaThru image formation model (Akkaynak and Treibitz 2019) within NeRF to successfully achieve novel view synthesis in real-world underwater scenes. ScatterNeRF (Ramazzina et al. 2023) uses inverse rendering, while DecoNeRF (Zhang et al. 2025) employs pseudo-labeling to disentangle scattering media from scene content and restore clear views in foggy scenes. Although these methods significantly improve reconstruction quality, they are often hindered by slow rendering speed and high memory consumption. Recently, 3DGS (Kerbl et al. 2023) has gained increasing attention for underwater scene reconstruction. SeaSplat (Yang et al. 2024a) incorporates underwater physics into 3D Gaussians to improve color restoration and structural fidelity, while WaterSplatting (Li et al. 2024a) combines implicit medium modeling with explicit object representation for efficient, high-quality reconstruction. UW-GS (Wang et al. 2024) further enhances geometry using pseudo-depth priors from foundation models (Yang et al. 2024b). Despite these advances, existing 3DGS-based methods have difficulty recovering accurate scene structure in underwater environments.
@@
Multi-view stereo (MVS) algorithms reconstruct detailed 3D geometry by leveraging images acquired from multiple calibrated viewpoints. Early approaches employed voxelization (Kutulakos and Seitz 2000), dense point clouds (Furukawa and Ponce 2009), and depth maps (Galliani et al. 2015; Schonberger and Frahm 2016). In the deep learning era, MVSNet (Yao et al. 2018) introduced cost volumes to form geometry-aware 3D representations, greatly enhancing accuracy. Building upon these advancements, MVS has been integrated into neural rendering (Mildenhall et al. 2020; Kerbl et al. 2023), where NeRF-based methods (Chen et al. 2021a; Wang et al. 2021; Fu et al. 2022; Wu et al. 2025) utilize these concepts for surface reconstruction, while 3DGS-based approaches (Chen et al. 2023; Guédon and Lepetit 2024; Huang et al. 2024; Chung et al. 2024; Chen et al. 2024a, b; Liu et al. 2024; Li et al. 2024b; Huang et al. 2025b) harness multi-view relationships to enhance geometric fidelity. Among these, recent efforts (Safadoust et al. 2024; Han et al. 2024) have demonstrated promising results by constructing virtual stereo pairs to enforce binocular stereo consistency, effectively regularizing 3D Gaussians.
@@
3D Gaussian Splatting (Kerbl et al. 2023) models a scene with anisotropic 3D Gaussian primitives, each parameterized by a center position $\mu\in\mathbb{R}^{3}$, covariance matrix $\Sigma\in\mathbb{R}^{3\times 3}$, color $c$, opacity $\alpha$, and view-dependent appearance modeled by spherical harmonics. Each 3D Gaussian defines its spatial influence at a point $\mathbf{X}\in\mathbb{R}^{3}$ centered at $\mu$ as:
@@
For rendering, 3D Gaussians are projected onto 2D screen-space and composited via alpha-blending of $N$ primitives that are sorted by depth and overlap at each pixel:
@@
To render a 3D underwater scene, the image formation is guided by the revised underwater image formation model (Akkaynak and Treibitz 2018), which decomposes the observed image into a direct component attenuated by the medium with respect to depth $z$, and a backscatter component accumulated along the viewing direction, as follows:
@@
Figure 2 shows the overall architecture of our method. We begin by initializing 3D Gaussians from a collection of RGB images $I\in\mathbb{R}^{H\times W\times 3}$, along with their corresponding camera intrinsics $K\in\mathbb{R}^{3\times 3}$ and extrinsic parameters $P\in\mathbb{R}^{4\times 4}$, which represent the homogeneous transformation matrix, recovered using Structure-from-Motion (SfM) (Schonberger and Frahm 2016). We model medium properties along the ray $\vec{\mathbf{r}}$ using an MLP $\phi_{\text{med}}$, following (Li et al. 2024a), which directly predicts attenuation $\sigma^{\text{attn}}$, backscattering $\sigma^{\text{bs}}$, and medium color $c^{\text{med}}$ as follows:
@@
We propose a trinocular stereo framework for 3DGS to improve object representation in scattering media. While prior works (Safadoust et al. 2024; Han et al. 2024) use only binocular stereo consistency, we show that a vertically translated viewpoint provides orthogonal constraints that complement horizontal stereo for 3D Gaussians. As shown in Figure 3, view-dependent sampling in alpha-blending often causes inconsistencies under minor viewpoint changes. Furthermore, previous stereo matching studies (Okutomi and Kanade 1993; Imran et al. 2020; Shamsafar and Zell 2021) have shown that multi-baseline stereo surpasses single-baseline methods in accuracy and robustness. Inspired by these observations, we enforce trinocular view consistency with orthogonal baselines of varying lengths to better regularize the spatial position of 3D Gaussians.
@@
From a given camera pose $P_{c}$, we obtain the rasterized object and medium components, denoted by $I^{\text{obj}}_{c}$ and $I^{\text{med}}_{c}$, respectively. We define two virtual viewpoints by translating $P_{c}$ along the horizontal $h$ and vertical $v$ axes as:
@@
where $\mathbb{I}$ is the $3\times 3$ identity matrix, $\mathbf{t}_{h}=(b_{h},0,0)^{\top}$ and $\mathbf{t}_{v}=(0,b_{v},0)^{\top}$ are translation vectors defining the horizontally and vertically translated camera poses $P_{h}$ and $P_{v}$, respectively, with $b_{h},b_{v}\in\mathbb{R}$ denoting the baseline distances. Using these poses, we render two images $I^{\text{obj}}_{h}$ and $I^{\text{obj}}_{v}$ from viewpoints $P_{h}$ and $P_{v}$. We then apply disparity-based inverse warping to the rendered images. Let $D_{c}\in\mathbb{R}^{H\times W\times 1}$ be the rendered depth map from $P_{c}$. Based on stereo geometry, the horizontal and vertical disparity maps are computed as:
@@
where $f_{h},f_{v}\in\mathbb{R}$ denote the focal lengths in the horizontal and vertical directions. The disparity maps $d_{h},d_{v}\in\mathbb{R}^{H\times W\times 1}$ are used to align the images rendered from the translated virtual viewpoints to the original view, and $(x,y)$ denotes the spatial pixel coordinates. Then, the rendered images from the virtual viewpoints, $I^{\text{obj}}_{h}$ and $I^{\text{obj}}_{v}$, are inverse warped using disparity maps $d_{h}$ and $d_{v}$, respectively, in the direction opposite to the virtual camera shift, aligning them with the central image rendered from $P_{c}$. The resulting inverse warped images $I^{\text{obj}}_{h\rightarrow c},I^{\text{obj}}_{v\rightarrow c}\in\mathbb{R}^{H\times W\times 3}$ are defined as:
@@
Afterward, to emphasize dark regions as perceived in human vision, we follow prior studies (Mildenhall et al. 2022; Levy et al. 2023; Li et al. 2024a), and employ a regularized L1 loss $L_{\text{R-L1}}$ defined as:
@@
To enhance geometric alignment, we introduce a synthetic epipolar depth prior $D_{\text{epi}}$, derived via triangulation between the translated viewpoints. For reliable correspondences in triangulation, we use 3D Gaussians sampled from the intersection of the view frusta of all trinocular camera views, with opacity above $\tau_{\alpha}$. Each selected Gaussian is projected onto the image planes of $P_{h}$ and $P_{v}$ via projection matrices $\mathbf{M}_{h},\mathbf{M}_{v}\in\mathbb{R}^{3\times 4}$, producing screen-space points $\mathbf{x}_{h}^{i},\mathbf{x}_{v}^{i}\in\mathbb{R}^{2}$. Let the homogeneous coordinates be $\tilde{\mathbf{x}}_{k}^{i}=[\mathbf{x}_{k}^{i\top},1]^{\top}\in\mathbb{R}^{3}$ and $\tilde{\mathbf{X}}_{i}=[\mathbf{X}_{i}^{\top},1]^{\top}\in\mathbb{R}^{4}$. We apply epipolar geometry to the projected correspondences, which yields a homogeneous linear system $\mathbf{A}_{i}\tilde{\mathbf{X}}_{i}=\mathbf{0}$, where $\mathbf{A}_{i}\in\mathbb{R}^{4\times 4}$ stacks the constraints from both views. To solve it stably, we rewrite it in the least-squares form:
@@
where $\mathbf{A}_{i}^{\prime}\in\mathbb{R}^{4\times 3}$ and $\mathbf{b}_{i}\in\mathbb{R}^{4}$ are obtained by separating the last column of $\mathbf{A}_{i}$. The triangulated point $\hat{\mathbf{X}}_{i}$ is transformed to the central camera frame via extrinsics $\mathbf{R}_{c}\in\mathbb{R}^{3\times 3}$, $\mathbf{t}_{c}\in\mathbb{R}^{3}$, and the depth prior is defined as the $z$-component:
@@
After that, we apply an edge-aware Log-L1 loss (Turkulainen et al. 2025) using $D_{\text{epi}}$ as follows:
@@
where $L_{\text{R-SSIM}}$ and $\lambda_{\text{s}}$ denote the Regularized SSIM loss (Li et al. 2024a) and its corresponding weight, respectively, with $\lambda_{\text{s}}=0.2$. In addition to the baseline, we incorporate our proposed components into the final training objective:
@@
We evaluate underwater scene reconstruction on four scenes from SeaThru-NeRF dataset (Levy et al. 2023) and two scenes from In-the-Wild dataset (Tang et al. 2024). In addition, to evaluate scene restoration performance, we generate simulated underwater and foggy scenes using the Fern scene from the LLFF dataset (Mildenhall et al. 2019). We simulate underwater scenes following previous methods (Levy et al. 2023) using normalized depth obtained from a depth foundation model (Yang et al. 2024b) with $\beta_{D}=[1.3,\ 1.2,\ 0.9]$, $\beta_{B}=[0.95,\ 0.85,\ 0.7]$, and $\beta_{\infty}=[0.07,\ 0.2,\ 0.39]$. For foggy scenes, we use $\beta_{D}=[0.5,\ 0.5,\ 0.5]$ and $\beta_{B}=\beta_{\infty}=1.2$. To validate our method, we compare scene quality using PSNR, SSIM (Wang et al. 2004), and LPIPS (Zhang et al. 2018).
@@
Our method builds on Nerfstudio (Tancik et al. 2023), and employs a TCNN-based MLP (Müller 2021). We use 7K/3K training steps for densification/finetuning on SeaThru-NeRF (Levy et al. 2023) and simulated datasets, and 10K/5K steps on In-the-Wild (Tang et al. 2024). At each iteration, $b_{v}$ is sampled from [0.4,0.4][-0.4,0.4], and $b_{h}$ is set to $1.5b_{v}$. We enable $L_{\text{tri}}$ and $L_{\text{epi}}$ during iterations 4K–8K and 6K–12K for 10K and 15K training steps, respectively. The $\tau_{\alpha}$ for synthetic epipolar depth is set to 0.8. The input resolution is increased from $\times 1/4$ to $\times 1/2$ at $1/5$ of total training steps, and from $\times 1/2$ to full resolution at $2/5$ to boost the training process. For depth-aware alpha adjustment, $w$ is set to 0.5 for $t<t_{\alpha}$ and is decayed to 0 at $t_{\alpha}$. We set the transition step $t_{\alpha}$ to 4K and 10K for 10K and 15K training steps, respectively. All experiments use a single RTX A6000.